OpenStudy (anonymous):

The length of 1 side of an equilateral triangle is 12. What is the distance from the centroid of the triangle to a side? (express in simplest radical form)

OpenStudy (anonymous):

answer is 6/ sqrt(3)

OpenStudy (anonymous):

\[\frac{6\sqrt{3}}{3}\]

OpenStudy (anonymous):

\[2 \sqrt{3}\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):